OpenStudy (anonymous):

use induction to prove that for all n > 0, n!>=n^(2-1), having trouble with the inductive step

OpenStudy (anonymous):

What is the equation? \[ n! \geq n^{2-1} \]?

OpenStudy (anonymous):

yes, not familiar with equation generator, sorry

OpenStudy (anonymous):

I mean isn't that the same as: \[ n! \geq n^1 \\ n! \geq n \]

OpenStudy (anonymous):

dang, sorry, wrote it wront, it's supposed to be n!>=2^(n-1)

OpenStudy (anonymous):

basis step is 0! = 1, then we assume true the hypothesis

OpenStudy (anonymous):

First plug in \(n+1\) into \(n\)

OpenStudy (anonymous):

(n+1)! >= 2^((n+1) - 1) >= 2^n

OpenStudy (anonymous):

Now you can pull out \(n+1\) and \(2\)?

OpenStudy (anonymous):

or the left side may be n(n-1)(n-2)...(1)(n+1)

OpenStudy (anonymous):

not sure what you mean pull out

OpenStudy (anonymous):

\((n+1)! = (n+1)n!\)

OpenStudy (anonymous):

yes, now i see, how does that relate to the right side?

OpenStudy (anonymous):

Now you can divide both sides by \(n+1\)

OpenStudy (anonymous):

\[ \large n! \geq \frac{2}{n+1}2^{n-1} \]

OpenStudy (anonymous):

Now you want to use the \[ n<0 \]inequality

OpenStudy (anonymous):

can you demonstrate that inequality?

OpenStudy (anonymous):

oops, I mean \(n > 0\)

OpenStudy (anonymous):

so you mean the original hypothesis now relates to this equation?

OpenStudy (anonymous):

yeah, it always did.

OpenStudy (anonymous):

remember that \[ n > 0 \implies n+1 > 1 \implies n+1 \geq 2 \implies \frac{2}{n+1} \geq 1 \]

OpenStudy (anonymous):

\[ n+1>2 \implies n+1 \geq 2 \]Comes from the fact that \(n\) can't be a non-integer.

OpenStudy (anonymous):

does that prove our hypothesis?

OpenStudy (anonymous):

now you combine the two equations we used.

OpenStudy (anonymous):

wio, can you help with this, is the inequality transitive, and can therefore be combined?

OpenStudy (anonymous):

\[ \frac{2}{n+1} \geq 1 \implies \frac{2}{n+1}2^{n-1} \geq 2^{n-1} \]

OpenStudy (anonymous):

thanks, gonna have to practice more of this to get it to sink in...

OpenStudy (anonymous):

Basically the steps are: 1) Plug in \(n+1\) 2) Try to get it in terms of \(n\) again.

OpenStudy (anonymous):

thanks again, signing off until tomorrow

OpenStudy (anonymous):

one question about this n>0⟹n+1>1⟹n+1≥2⟹2n+1≥1

OpenStudy (anonymous):

shouldn't the last part of it be 1>=2/(n+1)?